### v1

# What is the Right Notion of Distance between Predict-then-Optimize Tasks?

Abstract
Comparing datasets is a fundamental task in machine learning, essential for various learning paradigms—from evaluating train and test datasets for model generalization to using dataset similarity for detecting data drift. While traditional notions of dataset distances offer principled measures of similarity, their utility has largely been assessed through prediction error minimization. However, in Predict-then-Optimize (PtO) frameworks, where predictions serve as inputs for downstream optimization tasks, model performance is measured through decision regret minimization rather than prediction error minimization. In this work, we (i) show that traditional dataset distances, which rely solely on feature and label dimensions, lack informativeness in the PtO context, and (ii) propose a new dataset distance that incorporates the impacts of downstream decisions. Our results show that this decision-aware dataset distance effectively captures adaptation success in PtO contexts, providing a PtO adaptation bound in terms of dataset distance. Empirically, we show that our proposed distance measure accurately predicts transferability across three different PtO tasks from the literature.
1 Introduction
Comparing datasets is a fundamental task in machine learning and a crucial component of various downstream tasks. Understanding the similarity (or dissimilarity) of datasets can inform decisions in transfer learning (Tran et al., 2019; Ben-David et al., 2010), multitask learning (Janati et al., 2019; Shui et al., 2019), and data valuation (Just et al., 2023; Jiang et al., 2023), among other applications. For example, selecting a pre-training dataset that is similar to a data-poor target domain can lead to better fine-tuning performance. Notions of dataset distance have emerged as a principled way of quantifying these similarities and differences (Mercioni and Holban, 2019; Janati et al., 2019; Alvarez-Melis and Fusi, 2020). Such distances provide insights into the relation and correspondence between data distributions, help in evaluating model performance, and guide the selection of appropriate learning algorithms.
The concept of dataset can vary based on context and objectives. In classical statistics, it generally refers to feature vectors, focusing on the distribution and relationships within a feature space $\mathcal{X}$. Traditional statistical tests, such as the chi-squared test for categorical variables (Pearson, 1900) and the Kolmogorov-Smirnov test for numerical variables (Massey, 1951), quantify similarity based on features alone. Additionally, classic distributional distances offer formal measures of dataset similarity: the Total Variation distance (Verdú, 2014) quantifies the maximum discrepancy between distributions; Wasserstein distance, or Earth Mover’s Distance, measures the cost of transforming one distribution into another (Villani, 2008); and Integral Probability Metrics (IPM) measures how well a class of classifiers can distinguish samples from the two distributions (Müller, 1997).
In supervised learning, datasets include both features from space $\mathcal{X}$ and labels from space $\mathcal{Y}$. The distance between two such datasets involves measuring both the feature and label differences. This can be challenging when the label space $\mathcal{Y}$ is not a metric space. Approaches such as those proposed by Courty et al. (2014), Alvarez-Melis et al. (2018), and Alvarez-Melis and Fusi (2020) offer a principled method for computing dataset distances considering the joint feature-label distribution $\mathcal{P}(\mathcal{X}\times\mathcal{Y})$. These methods ensure that both the features and labels are adequately accounted for in the distance measure, offering a more holistic comparison between datasets.
However, the Predict-then-Optimize (PtO) framework introduces a unique challenge by using machine learning predictions as inputs for a downstream optimization problem, shifting the focus from minimizing prediction error to minimizing decision regret (Agrawal et al., 2019; Amos and Kolter, 2017; Elmachtoub and Grigas, 2022; Mandi et al., 2023). This results in PtO tasks involving not just a feature-label dataset, but also a decision space $\Omega$ of optimization solutions, creating a feature-label-decision dataset with samples in $\mathcal{X}\times\mathcal{Y}\times\Omega$. The decision space $\Omega$ may not be a metric space; for example, decisions related to the solution of a top-k problem do not necessarily form a metric space. Moreover, decisions might need to be evaluated under various criteria, such as minimizing travel distance versus maximizing safety. Even if $\Omega$ were a metric space, it is uncertain whether its associated distance would be meaningful for assessing the adaptability of a PtO task across different domains. This complexity underscores the need for new distance measures that incorporate decisions to accurately capture the nature of PtO tasks.
In this work, we introduce a decision-aware dataset distance measure based on Optimal Transport (OT) techniques (Villani, 2008) that incorporates features, labels, and decisions. Our decision-aware dataset distance is the first to integrate and capture downstream decisions as part of the dataset, addressing the unique challenges of PtO tasks. To evaluate its effectiveness, we use this distance as a learning-free criterion for assessing transferability under distribution shifts. In the context of domain adaptation, where PtO success is determined by minimizing decision regret rather than prediction error (Elmachtoub and Grigas, 2022; Mandi et al., 2023), we derive a domain adaptation bound that emphasizes the necessity of incorporating features, labels, and decisions together. Our empirical analysis spans three PtO tasks from the literature—Linear Model Top-K, Warcraft Shortest Path, and Inventory Stock Problem—demonstrating that our decision-aware distance better predicts transfer performance compared to feature-label distances alone.
2 Related Work
Dataset Distances via Optimal Transport
Optimal Transport (OT)-based distances have gained traction as an effective method for comparing datasets. These methods characterize datasets as empirical probability distributions supported in finite samples, and require a cost function between pairs of samples to be provided as an input. Most OT-based dataset distance approaches define this cost function solely in terms of the features of the data, either directly or in a latent embedding space. For example, Muzellec and Cuturi (2018) proposed representing objects as elliptical distributions and scaling these computations, while Frogner et al. (2019) extended this to discrete measures. Delon and Desolneux (2020) introduced a Wasserstein-type distance for Gaussian mixture models. These approaches are useful mostly in unsupervised learning settings since they do not take into account labels or classes associated with data points. To address this limitation, a different line of work has proposed extensions of OT amenable to supervised or semi-supervised learning settings that explicitly incorporate label information in the cost function. Courty et al. (2014) used group-norm penalties to guide OT towards class-coherent matches while Alvarez-Melis et al. (2018) employed submodular cost functions to integrate label information into the OT objective. For discrete labels, Alvarez-Melis and Fusi (2020) proposed using a hierarchical OT approach to compute label-to-label distances as distances between the conditional distributions of features defined by the labels.
Predict-then-Optimize (PtO)
In recent years, the PtO framework has advanced significantly in integrating machine learning with downstream optimization. The frameworks proposed by Amos and Kolter (2017); Donti et al. (2017); Wilder et al. (2018); Elmachtoub and Grigas (2022) have been instrumental in this integration. Subsequent contributions have focused on differentiating through the parameters of optimization problems with various structures, including learning appropriate loss functions (Wang et al., 2020; Shah et al., 2022, 2023; Bansal et al., 2023) and handling nonlinear objectives (Qi et al., 2023; Elmachtoub et al., 2023). Recently, attention has also been given to data-driven challenges within the PtO framework, including robustness to adversarial label drift (Johnson-Yu et al., 2023) and active learning for data acquisition (Liu et al., 2023). These works introduce learning-based algorithms tailored to specific challenges, yet they converge on a common underlying paradigm: dataset similarity. For label drift, the issue centers on the (dis)similarity between training and test datasets. In data acquisition, it involves assessing the (dis)similarity, in terms of diversity and informativeness (Cacciarelli and Kulahci, 2023), between the training dataset and the acquisition source.
3 Background
3.1 Optimal Transport
OT theory provides an elegant and powerful mathematical framework for measuring the distance between probability distributions by characterizing similarity in terms of correspondence and transfer (Villani, 2008; Kantorovitch, 1942). In a nutshell, OT addresses the problem of transferring probability mass from one distribution to another while minimizing a cost function associated with the transportation.
Formally, given two probability distributions $\alpha$ and $\beta$ defined on measurable spaces $\mathcal{X}$ and $\mathcal{Y}$, respectively, the OT problem seeks a transport plan $\pi$ (defined as a coupling between $\alpha$ and $\beta$) that minimizes the total transportation cost. According to the Kantorovich formulation (Kantorovitch, 1942), for any coupling $\pi$, the transport cost between $\alpha$ and $\beta$ with respect to $\pi$ is defined as:
$d_{T}(\alpha,\beta;\pi):=\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)\mathop{}\!\mathrm{d}\pi(x,y)$ (1)
where $c(x,y)$ is the cost function representing the cost of transporting mass from point $x\in\mathcal{X}$ to point $y\in\mathcal{Y}$. The transport cost $d_{T}(\alpha,\beta;\pi)$ defines a distance, known as the transport distance with respect to $\pi$, between $\alpha$ and $\beta$. The OT problem then minimizes the transport cost over all possible couplings between $\alpha$ and $\beta$, defining the optimal transport distance as:
$d_{OT}(\alpha,\beta):=\min\limits_{\pi\in\Pi(\alpha,\beta)}\int_{\mathcal{X}\times\mathcal{Y}}c(x,y)\mathop{}\!\mathrm{d}\pi(x,y)$ (2)
where $\Pi(\alpha,\beta)$ denotes the set of all possible couplings (transport plans) that have $\alpha$ and $\beta$ as their marginals. This formulation finds the optimal way to transform one distribution into another by minimizing the total transportation cost.
3.2 Dataset Distances via Optimal Transport
In supervised machine learning, datasets can be represented as empirical joint distributions over a feature-label space $\mathcal{X}\times\mathcal{Y}$. OT distances can be used to measure the similarity between these empirical distributions, thus providing a principled way to compare datasets. Given two datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ consisting of feature-label tuples $(x,y)$ and $(x^{\prime},y^{\prime})$, respectively, the challenge of defining a transport distance between $\mathcal{D}$ and $\mathcal{D}^{\prime}$ lies in the challenge of defining an appropriate cost function between $(x,y)$ and $(x^{\prime},y^{\prime})$ pairs. A straightforward way to define the feature-label pairwise cost is via the individual metrics in $\mathcal{X}$ and $\mathcal{Y}$ if available. If $d_{\mathcal{X}}$ and $d_{\mathcal{Y}}$ are metrics on $\mathcal{X}$ and $\mathcal{Y}$, respectively, the cost function can be defined as:
$c((x,y),(x^{\prime},y^{\prime}))=\big{(}d_{\mathcal{X}}(x,x^{\prime})^{p}+d_{\mathcal{Y}}(y,y^{\prime})^{p}\big{)}^{1/p}$
for $p\geq 1$. This point-wise cost function defines a valid metric on $\mathcal{X}\times\mathcal{Y}$. However, it is uncommon for $d_{\mathcal{Y}}$ to be readily available. To address this, Courty et al. (2017) propose replacing $d_{\mathcal{Y}}(y,y^{\prime})$ with a loss function $\mathcal{L}(y,y^{\prime})$ that measures the discrepancy between $y$ and $y^{\prime}$ while Alvarez-Melis and Fusi (2020) suggest using a p-Wasserstein distance between the conditional distributions of features defined by $y$ and $y^{\prime}$ as an alternative to $d_{\mathcal{Y}}(y,y^{\prime})$.
3.3 Predict-then-Optimize
The Predict-then-Optimize (PtO) framework involves two sequential steps: prediction and optimization. First, a predictive model $f$ is used to predict costs based on some features $x_{1},\ldots,x_{N}\in\mathcal{X}$, represented as ${\bf\it\hat{y}}=[\hat{y}_{1},\ldots,\hat{y}_{N}]=[f(x_{1}),\ldots,f(x_{N})]$. Second, an optimization model uses these predicted costs $^{\bf\it\hat{y}}$ as the objective function costs:
$M({\bf\it\hat{y}}):=\text{argmax}_{w}\ g(w;{\bf\it\hat{y}}),\hskip 11.38092pts.t.\hskip 5.69046ptw\in\Omega$ (3)
where $\Omega$ is the space of feasible solutions. We assume that $w^{*}:\mathbb{R}^{d}\to\Omega$ acts as an oracle for solving this optimization problem, such that $w^{*}({\bf\it\hat{y}})$ represents the optimal solution for $M({\bf\it\hat{y}})$. However, the solution $w^{*}({\bf\it\hat{y}})$ is optimal for $M({\bf\it\hat{y}})$ but might not be optimal for $M({\bf\it y})$, where ${\bf\it y}$ represents the true costs.
Given a hypothesis function $f:\mathcal{X}\to\mathcal{Y}$, we measure its performance on the optimization problem $M({\bf\it y})$ using the predicted cost vector ${\bf\it\hat{y}}=[f(x_{1}),\ldots,f(x_{N})]$ and the true cost vector ${\bf\it y}=[y_{1},\ldots,y_{N}]$. This is quantified as the decision quality $q(\mathbf{\hat{y}},{\bf\it y})=g(w^{*}(\mathbf{\hat{y}});{\bf\it y})$, reflecting the quality of decisions made using $w^{*}(\mathbf{\hat{y}})$ as a solution to $M({\bf\it y})$. The decision quality regret, which evaluates the performance of $f$, is defined as:
$q_{\text{reg}}({\bf\it\hat{y}},{\bf\it y})=|q({\bf\it y},{\bf\it y})-q({\bf\it\hat{y}},{\bf\it y})|$ (4)
The goal of decision-focused learning in a PtO task is to learn a predictive model $f_{\theta}$ that minimizes the decision quality regret, ensuring that the decisions derived from the predictions are as close to optimal as possible.
4 Decision-aware Dataset Distances
Our primary objective is to develop a formal notion of distance between PtO tasks. Specifically, we aim to define a distance $d(\mathcal{D},\mathcal{D}^{\prime})$ between datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ that reflects task similarity in the context of PtO, and hence, is predictive of decision regret transferability. To achieve this, we consider datasets $\mathcal{D}$ and $\mathcal{D}^{\prime}$ consisting of feature-label-decision tuples $(x,y,z)\in\mathcal{X}\times\mathcal{Y}\times\Omega$, where $\mathcal{X}$ is the feature space, $\mathcal{Y}$ the label space, and $\Omega$ the decision space. For a decision space defined by a downstream optimization problem $M(\cdot)$ (Eq. 3), the decision space $\Omega$ can be considered as the set of feasible solutions of the optimization problem $M(y)$ for all $y\in\mathcal{Y}$. We will refer to the joint feature-label-decision space as $\mathcal{W}:=\mathcal{X}\times\mathcal{Y}\times\Omega$.
Given two distributions $\mathcal{P}$ and $\mathcal{P}^{\prime}$ over $\mathcal{W}$, we consider two datasets drawn from these distributions: $\mathcal{D}=\{(x_{i},y_{i},z_{i})\}_{i=1}^{n}\sim\mathcal{P}$ and $\mathcal{D}^{\prime}=\{(x^{\prime}_{i},y^{\prime}_{i},z^{\prime}_{i})\}_{i=1}^{m}\sim\mathcal{P}^{\prime}$. Our objective is to ensure that the distance $d(\mathcal{D},\mathcal{D}^{\prime})$ accurately predicts transferability between $\mathcal{P}$ and $\mathcal{P}^{\prime}$, with transferability assessed in terms of decision regret minimization.
Motivating Example
To highlight the importance of incorporating decisions, in addition to features and labels, when comparing PtO tasks, we evaluate task performance and similarity under distribution shifts in the Linear Model Top-K setting from Shah et al. (2022). The task is as follows:
  • Predict: Given the feature $x_{n}\sim\mathcal{P}_{\mathcal{X}}$, where $\mathcal{P}_{\mathcal{X}}=\text{Unif}[-1,1]$, of a resource $n$, use a linear model to predict its utility $\hat{y}_{n}$, where the true utility $y_{n}=p(x_{n})$ is a cubic polynomial in $x_{n}$. Combining predictions for $N$ resources yields $\hat{\bm{y}}=[\hat{y}_{1},\ldots,\hat{y}_{N}]$.
  • Optimize: Choose $K=1$ out of $N$ resources with the highest utility. This corresponds to solving the optimization problem $M(\hat{\bm{y}})=\max_{\bm{z}\in[0,1]^{N}}\{\bm{z}\cdot\sigma_{x}(\hat{\bm{y}})\}$ such that $||\bm{z}||_{0}=K$, where $\sigma_{x}$ is the permutation that orders $\hat{\bm{y}}$ in ascending order of $\bm{x}=[x_{1},\ldots,x_{N}]$.
Consider the distributions with target shifts, $Z\gamma:=[\text{id},p_{\gamma},w^{*}\circ p_{\gamma}]_{*}\mathcal{P}_{\mathcal{X}}$, denoted as the pushforward distribution of $\mathcal{P}_{\mathcal{X}}$ through the composite mapping $[\text{id},p_{\gamma},w^{*}\circ p_{\gamma}]$. Here, id is the identity function, $p_{\gamma}=10(x^{3}-\gamma x)$ is the labeling function with target shift parameter $\gamma$, and $w^{*}\circ p_{\gamma}$ denotes the composition of the optimization oracle $w^{*}$ with $p_{\gamma}$. Let $Z_{0}$ be source distribution A, $Z_{1.2}$ be source distribution B, and $Z_{0.65}$ be the target distribution (as in Shah et al. (2022)). Figure 1 illustrates an instance drawn from each of these domains.
Figure 1: Linear Model Top-K instances under target shift
Given the optimal weights $\theta_{A}$ and $\theta_{B}$ of a linear model $f(\cdot;\theta)$, learned in a decision-focused manner using training datasets $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$ from source distributions A and B, we need to perform Top-K selection on a dataset $\mathcal{D}_{C}$ drawn from the target distribution $Z_{0.65}$. Assume we lack the resources to train new weights on $\mathcal{D}_{C}$—perhaps having only a single optimization instance, insufficient for learning new weights. Thus, we must choose between $\theta_{A}$ and $\theta_{B}$ for the PtO task on $\mathcal{D}_{C}$.
A potential approach is to use dataset distance as a criterion to select the source distribution most similar to the target distribution. However, if we rely on traditional OT dataset distances, which consider only features and labels, we would be indifferent between $\theta_{A}$ and $\theta_{B}$ since $d_{OT}(\mathcal{D}_{A},\mathcal{D}_{C})=d_{OT}(\mathcal{D}_{B},\mathcal{D}_{C})$ (Appendix Fig. 8). Yet, $\theta_{A}$ leads to zero regret, while $\theta_{B}$ results in a regret close to 4, making $\theta_{A}$ clearly more suitable for the PtO task on $\mathcal{D}_{C}$.
This difference in performance, illustrated in Figure 1 by the Top-K decisions derived from each model, suggests that dataset distances should reflect the significant disparity in regret, indicating that $\mathcal{D}_{C}$ is closer to $\mathcal{D}_{A}$ than to $\mathcal{D}_{B}$. Therefore, we argue that feature-label distances alone are insufficient, and incorporating decision components is necessary for distances to accurately reflect adaptability in PtO tasks.
4.1 Incorporating Decision into OT Distances
A dataset for a PtO task consists of feature-label-decision triplets $(x,y,z)\in\mathcal{X}\times\mathcal{Y}\times\Omega$. To compute an OT-based dataset distance (as defined in Section 3.2) between PtO datasets, we need a suitable metric for the joint space $Z:=\mathcal{X}\times\mathcal{Y}\times\Omega$ to serve as the ground cost function in the OT problem.
Pairwise distances between feature-label samples can be computed using metrics on the feature space $\mathcal{X}$ and the label space $\mathcal{Y}$. If $\mathcal{Y}$ lacks a natural metric, we can use the metric on $\mathcal{X}$ to compare labels based on their association with features (Alvarez-Melis and Fusi, 2020). Defining a metric for the decision space $\Omega$ is also challenging, as it may not inherently form a metric space. For example, decisions related to resource allocation or scheduling may not align with traditional metrics. Even when $\Omega$ is equipped with a metric, it might not reflect decision quality regret effectively—a key aspect for PtO tasks. For instance, in comparing paths in a $p\times q$ grid, Euclidean or Manhattan distances will not capture differences under different goals like minimizing cost versus maximizing safety.
Therefore, the metric in $Z$ must accurately reflect decision quality regret, as outlined in Equation 4. This means the metric should not only measure distances between feature-label-decision triplets but also incorporate the quality and impact of decisions. To facilitate this, we define the decision quality disparity to compare decisions $z$ and $z^{\prime}$ in $\Omega$ with respect to their quality under a pair of labels $y$ and $y^{\prime}$:
Definition 4.1 (Decision quality disparity).
For an optimization problem $M$ with objective function $g$, the decision quality disparity function $l_{g}(\ \cdot\ ;y,y^{\prime}):\Omega^{2}\to\mathbb{R}$ measures the difference in decision quality between two decisions $z,z^{\prime}\in\Omega$ given the labels $y,y^{\prime}\in\mathcal{Y}$. It is defined as:
$l_{g}(z,z^{\prime};y,y^{\prime}):=\absolutevalue{g(z;y)-g(z^{\prime};y^{\prime})}$ | (5)
Note that decision quality regret (defined in Section 3.3) is a special case of decision quality disparity, where $q_{\text{reg}}(\hat{y},y)=l_{g}(w^{*}(\hat{y}),w^{*}(y);y,y)$ for an optimizatin oracle $w^{*}$. We use this measure of difference led by decisions to define a point-wise notion of distance in $\mathcal{X}\times\mathcal{Y}\times\Omega$ that accounts for differences in features, labels, and decisions:
$c_{PtO}^{\bm{\alpha}}((x,y,z),(x^{\prime},y^{\prime},z^{\prime}))\triangleq$ $\ \ \alpha_{X}\cdot d_{\mathcal{X}}(x,x^{\prime})$
$+\alpha_{Y}\cdot d_{\mathcal{Y}}(y,y^{\prime})$ (6)
$+\alpha_{W}\cdot l_{g}(z,z^{\prime};y^{\prime},y^{\prime})$
[…]
Proposition 4.2.
For any $\bm{\alpha}=(\alpha_{X},\alpha_{Y},\alpha_{W})$ such that $\alpha_{X},\alpha_{Y},\alpha_{W}>0$, the decision-aware dataset distance $d_{OT}(\mathcal{D},\mathcal{D}^{\prime};c^{\bm{\alpha}}_{PtO})$ defines a valid metric on $\mathcal{P}(\mathcal{X}\times\mathcal{Y}\times\Omega)$, where $\mathcal{P}(\mathcal{X}\times\mathcal{Y}\times\Omega)$ represents the space of probability measures over joint distributions of features $\mathcal{X}$, labels $\mathcal{Y}$, and decisions $\Omega$. If $\alpha_{Y}=0$, then the decision-aware dataset distance is still at least a pseudometric.
This decision-aware dataset distance compares decisions $z$ and $z^{\prime}$ by evaluating their decision quality disparity in $\mathbb{R}$ relative to a pair of fixed labels, rather than directly comparing them in the decision space $\Omega$. Intuitively, comparing decisions based on their quality, i.e., comparing $g(z;y)$ with $g(z^{\prime};y)$, rather than comparing $z$ and $z^{\prime}$ directly using some metric in $\Omega$, if available, is reasonable because similar decisions might yield significantly different outcomes in the objective function. Our next result shows that this approach provides a reasonable measure for assessing adaptation success of PtO tasks between two distributions in the feature-label-decision space.
4.2 Adaptation Bound
Given source and target distributions $\mathcal{P}_{S}$ and $\mathcal{P}_{T}$ over $\mathcal{X}\times\mathcal{Y}$, along with a downstream optimization problem $M(\cdot)$ and a corresponding optimization oracle $w^{*}$, we tackle the domain adaptation problem within the PtO framework. Our goal is to bound the expected decision quality regret of a labeling function $f$ on the target domain $\mathcal{P}_{T}$, based on its distance to the source distribution $\mathcal{P}_{S}$. We achieve this by leveraging the decision-aware distance introduced in Section 4.1.
Previous work has bounded target error using expressions where the adjustable terms, minimized to achieve tighter bounds, correspond to dataset distances considering only features and labels (Courty et al., 2017). These types of bounds, with dataset distance terms, have been applied to loss functions that are bounded, symmetric, $k$-Lipschitz, and satisfy the triangle inequality. However, in PtO tasks, the error is measured as decision quality regret $q_{\text{reg}}$, which is inherently non-symmetric.
To address this, we use our proposed measure, the decision quality disparity $l_{q}$, to bound $q_{\text{reg}}$ by fixing the labels against which decision quality is assessed. Then, under Assumption 4.3, which ensures that the decision quality function has a bounded rate of change with respect to both the predicted and true cost vectors, provide a bound for the expected decision quality regret in the target distribution by using our decision-aware dataset distance (Theorem 4.4). This bound, expressed in terms of a feature-label-decision distance, underscores the importance of incorporating decision components into dataset distances. Highlighting that decision-aware distances more accurately reflect transferability in terms of decision regret, rather than those relying solely on feature-label components.
Let $f:\mathcal{X}\to\mathcal{Y}$ be any labeling function. We define the expected cost of $f$ under a distribution $\mathcal{P}$ over $\mathcal{X}\times\mathcal{Y}$ with respect to any cost function $l:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}$ as
$err(f;l,\mathcal{P}):=\mathbb{E}_{(x,y)\sim\mathcal{P}}\ l(f(x),y).$ (7)
Then, the expected decision quality regret of a labeling function $f$ under a distribution $\mathcal{P}$ is given by
$err(f;q_{\text{reg}},\mathcal{P})$ $=\mathbb{E}_{(x,y)\sim\mathcal{P}}\ q_{\text{reg}}(f(x),y)$
For any labeling function $f$, we provide a bound on the target decision quality regret $err(f;q_{\text{reg}},\mathcal{P}_{T})$. We provide this bound under Assumption 4.3 which ensures that the decision quality function has a bounded rate of change with respect to both the predicted and true cost vectors. As we highlight in lemmas B.2 and B.3 in the supplementary material, this is a reasonable assumption for common PtO tasks.
Assumption 4.3.
The decision quality function is $k_{1},k_{2}$-Lipschitz. This means that for any $y,y^{*},z,z^{*}\in\mathcal{Y}$ the following inequality holds:
$|q(y,y^{*})-q(z,z^{*})|\leq k_{1}\norm{y-z}+k_{2}\norm{y^{*}-z^{*}}$
Theorem 4.4.
Suppose Assumption 4.3 holds. Define the distributions $\mathcal{P}_{T}^{f}:=(x,y,w^{*}(f(x)))_{(x,y)\sim\mathcal{P}_{T}}$ and $\mathcal{P}_{S}^{*}:=(x,y,w^{*}(y))_{(x,y)\sim\mathcal{P}_{S}}$ over the joint feature-label-decision space $\mathcal{W}$. Let $\Pi^{*}$ be the coupling that minimizes the OT problem with ground cost $c_{PtO}^{\bm{\alpha}}$ between $\mathcal{P}_{T}^{f}$ and $\mathcal{P}_{S}^{*}$. Let $\tilde{f}$ be a labeling function that is $\phi$-Lipschitz transferable w.r.t. $\Pi^{*}$ [^1] This is a probabilistic notion of Lipschitz transferability between two distributions with a coupling $\Pi^{*}$ introduced by Courty et al. (2017). See definition in the supplementary material.. We assume $\mathcal{X}$ is bounded by $K$ and $\tilde{f}$ is $l$-Lipschitz, such that $|\tilde{f}(x_{1})-\tilde{f}(x_{2})|\leq 2lK=L$. Then, for all $\lambda>0$ and $\alpha_{W}\in(0,1)$ such that $(\lambda k_{1}+k_{2}+1)\alpha_{W}=1$, and $\alpha_{X}=\lambda k_{1}\alpha_{W}$ and $\alpha_{Y}=k_{2}\alpha_{W}$, we have with probability at least $1-\delta$ that:
$err(f;q_{\text{reg}},\mathcal{P}_{T})\leq$ $err(\tilde{f};q_{\text{reg}},\mathcal{P}_{S})+err(\tilde{f};q_{\text{reg}},\mathcal{P}_{T})$
$+k_{1}L\phi(\lambda)$
$+\frac{1}{\alpha_{W}}d_{OT}(\mathcal{P}_{T}^{f},\mathcal{P}_{S}^{*}\ ;c_{PtO}^{\bm{\alpha}})$
The proof of Theorem 4.4 is provided in the supplementary material. The first two terms correspond to the joint decision quality regret minimizer between the source and target distributions, indicating that domain adaptation in the PtO framework is effective only if we can attain predictions that lead to low decision quality regret in both domains simultaneously. This is similar to results in the literature of domain adaptation for supervised and unsupervised learning (Courty et al., 2017; Mansour et al., 2023; Ben-David et al., 2010). The third term $k_{1}L\phi(\lambda)$ measures the extend to which Lipschitzness between the source and target distributions does not hold.
The remaining term quantifies the discrepancy between the source domain $\mathcal{P}^{*}_{S}$ and the predicted target domain $\mathcal{P}^{f}_{T}$ using the optimal transport distance between their joint distributions of features, labels, and decisions. The bound depends on two parameters, $\lambda$ and $\alpha_{W}$: $\lambda$ controls the Lipschitz term and is valid for any $\lambda>0$, while $\alpha_{W}$ defines the weight assigned to decisions in the convex combination $c_{PtO}^{\bm{\alpha}}$. Note that the bound is valid for any combination of weights $\alpha_{X},\alpha_{Y},\alpha_{W}$, as $\lambda$ can always be chosen to result in a fixed convex combination. This contrasts with traditional domain adaptation methods, which rely on distances based solely on features and labels (Courty et al., 2017). Our approach recognizes the necessity of incorporating decisions into dataset distances when used for domain adaptation purposes for PtO tasks. Our OT-based dataset distance, defined by the ground cost function $c_{PtO}^{\bm{\alpha}}$, jointly accounts for differences in all key components—features, labels, and decisions—providing a comprehensive measure that is meaningful for adaptability of PtO tasks.
4.3 Weighting Ground Cost Components
In our approach, the weights on the ground cost components, $\bm{\alpha}$ in Eq. 4.1, are pivotal in defining the dataset distance, offering a flexible framework to account for the varying importance of features, labels, and decisions in PtO tasks. Unlike previous OT-based dataset distances that did not differentiate between the weights of feature and label components in the ground cost function—often because both were measured in the same space (Alvarez-Melis and Fusi, 2020) or were weighted equally (Courty et al., 2017)— our method allows for distinct weights, enabling a more nuanced evaluation of dataset similarity tailored to each specific task. This flexibility ensures that the distance metric reflects the relative significance of each dataset component according to its impact on the PtO task, which can vary widely in practice depending on the application. Indeed, our experiments show that different choices of weights can lead to significant changes in the behavior of the dataset distance in terms of capturing task-specific similarities and differences.
5 Experiments
In this section, we compare OT-dataset distances based on feature-label dimensions to our decision-aware OT-dataset distance, which also includes decision dimensions. A primary motivation for introducing a dataset distance that is informative of PtO performance was to offer a learning-free criterion for selecting a source datasets to train model on. We evaluate the extent to which these distances effectively compare PtO tasks in terms of their transferability.
(a) Linear Model Top-K
(b) Warcraft Shortest Path
(c) Inventory Stock
Figure 2: Effect of weights on predicting transferability. The color scale shows the R-squared value from a linear regression of regret transferability against dataset distance for various weights on features, labels, and decisions.
5.1 Experimental settings
We conduct our experiments using three predict-then-optimize tasks, chosen for their diverse domains and varying sensitivity to distribution shifts, making them ideal for analyzing dataset distances in domain adaptation contexts.
Linear Model Top-K
This setting, first proposed by Shah et al. (2022), involves training a linear model to map features $x_{n}\sim U[-1,1]$ to true utilities based on a cubic polynomial $p(x_{n})=10(x_{n}^{3}-0.65x_{n})$. The downstream task involves selecting the $K$ elements with the highest utility from the predicted values. We introduce synthetic distribution shifts by modifying the original feature-label distribution $\mathcal{P}=(\text{Id},p)_{*}U[-1,1]$. Specifically, for various values of $\gamma\in[0,1.3]$, we define the feature-label distributions $\mathcal{P}_{\gamma}=(\text{Id},p_{\gamma})_{*}U[-1,1]$ where $p_{\gamma}(x_{n})=10(x_{n}^{3}-\gamma x_{n})$, using $\mathcal{P}_{0.65}$ as the target distribution.
Warcraft Shortest Path
Adapted from Mandi et al. (2023) and Tang and Khalil (2023), this setting involves finding the minimum cost path on $d\times d$ RGB grid maps from the Warcraft II tileset dataset (Vlastelica et al., 2020), where each pixel has an unknown cost. The task is to predict these costs and then determine the minimum cost path from the top-left to the bottom-right pixel. The original distribution $\mathcal{P}$, which we treat as the target distribution, is defined over $\mathbb{R}^{d\times d}\times\mathbb{R}^{p\times p}$, where $d=96$ and $p=12$. Here, $\mathbb{R}^{d\times d}$ represents the feature space depicting maps, while $\mathbb{R}^{p\times p}$ represents the traveling costs on these maps. We induce a target shift for $\mathcal{P}_{\gamma}$ by uniformly sampling the costs for different pixel classes from the same range as $\mathcal{P}$ ([0.8, 9.2] for the Warcraft II tileset dataset).
Inventory Stock Problem
Adapted from Donti et al. (2017), this problem involves determining the optimal order quantity $z$ to minimize costs given a stochastic demand $y$, influenced by features $x$. The cost structure $f_{\rm stock}$ includes linear and quadratic costs for both ordering and deviations (over-orders and under-orders) from demand. Given a probability model $p(y|x;\theta)$, the optimization problem is: $\min_{z}\mathbb{E}_{y\sim p(y|x;\theta)}\left[f_{\text{stock}}(y,z)\right]$. Assuming discrete demands $d_{1},\ldots,d_{k}$ with probabilities $p\left(y=d_{i}|x;\theta\right)$, the problem can be expressed as a joint quadratic program (see supplementary material for details). We generate problem instances by randomly sampling $x\in\mathbb{R}^{n}$ and then generating $p(y|x;\theta)$ according to $p(y|x;\theta)\propto\exp((\Theta^{T}x)^{2})$. We introduce distribution shifts for both feature $x$ and label $y$. Specifically, $x$ is sampled from a Gaussian distribution where the mean is sampled from $U[-0.5,0.5]$, and $\Theta$ is also sampled from a Gaussian distribution.
5.2 Predicting Domain Transferability
To evaluate how effectively our decision-aware dataset distance measures transferability between PtO tasks, we analyze the correlation between the distance from a source dataset $\mathcal{D}_{S}$ to a target dataset $\mathcal{D}_{T}$ and the regret achieved on $\mathcal{D}_{T}$ by a model trained on $\mathcal{D}_{S}$. We assess regret transferability by determining how well the dataset distance predicts the regret performance on $\mathcal{D}_{T}$ based on the model trained on $\mathcal{D}_{S}$. Specifically, regret transferability is computed as:
$\mathcal{T}(\mathcal{D}_{S}\to\mathcal{D}_{T})=\frac{\text{regret}(\theta_{T},\mathcal{D}_{T})-\text{regret}(\theta_{S},\mathcal{D}_{T})}{\text{regret}(\theta_{T},\mathcal{D}_{T})}$
where $\theta_{T}$ and $\theta_{S}$ are the model weights learned by minimizing decision quality regret on datasets $\mathcal{D}_{T}$ and $\mathcal{D}_{S}$, respectively, and $\text{regret}(\theta,\mathcal{D})$ denotes the mean regret achieved by a model with weights $\theta$ over the optimization instances in dataset $\mathcal{D}$.
For each experimental setting described in Section 5.1, we consider $k$ source datasets $\mathcal{D}_{S_{1}},\ldots,\mathcal{D}_{S_{k}}$ sampled from different distributions and a target dataset $\mathcal{D}_{T}$, sampled from a target distribution. For each pair of source and target dataset $(\mathcal{D}_{S_{i}},\mathcal{D}_{T})$, we compute the regret transferability $\mathcal{T}(\mathcal{D}_{S_{i}}\to\mathcal{D}_{T})$  [^2] To compute $\mathcal{T}(\mathcal{D}_{S_{i}}\to\mathcal{D}_{T})$, we train model weights $\theta_{S_{i}}$ on $\mathcal{D}_{S_{i}}$ and $\theta_{T}$ on $\mathcal{D}_{T}$ using established decision-focused learning approaches. See supplementary material, Section D, for details. and analyze it in relation to our decision-aware dataset distance.
Figure 2 illustrates the correlation strength (measured by the R-squared value from linear regression) between regret transfer and dataset distance for different weighting combinations $\bm{\alpha}$. The analysis reveals that, in the Linear Model Top-K and Warcraft settings, incorporating the decision component ($\alpha_{W}>0$) significantly enhances the correlation between dataset distance and regret transfer, even when the label component ($\alpha_{Y}=0$) is excluded. Conversely, omitting the decision component ($\alpha_{W}=0$, shown on the left side of the plot) weakens this correlation. This trend is further highlighted in with optimal weighting combination such that $\alpha_{W}>0$ (Figure 3 [right]) is far more predictive of regret transfer than combinations that consider only features and labels (Figure 3 [left]).
Figure 3: Distance vs Adaptation in the Warcraft setting for two different ground cost weighing profiles (Sec. 4.3).
The advantage of including the decision component over the label component is less pronounced in the Inventory Stock problem (Fig. 2(c)). Here, either the label or decision component with features still maintains a strong correlation between regret transfer and dataset distance. To explore this further, we examine how differences in the label space $d_{y}(y,y^{\prime})$ correlate with differences in the decision space $l_{q}(y,y^{\prime},z,z^{\prime})$. In the inventory stock problem, there is a strong correlation between these differences (Fig. 4(b)), suggesting that decisions are closely tied to the labels. In contrast, the Warcraft domain lacks this strong correlation (Fig. 4(a)), making the decision component more critical for accurately predicting transferability.
(a) Warcraft Shortest Path
(b) Inventory Stock
Figure 4: Difference in labels against difference in decisions
5.3 Impact of Target Shift on PtO Similarity
Target shift, where target label distributions change while feature distributions remain constant, challenges domain adaptation in supervised learning by causing mismatches between source and target domains, often leading to poor transferability (Japkowicz and Stephen, 2002; Zhang et al., 2013). However, out experimental results show that some source datasets with high feature-label distance due to significant target shift, exhibit low PtO transferability (Fig.3 [left]). This suggests that the impact of target shift may not directly correlate with PtO transferability (Fig.4(a)), and can be less pronounced in PtO contexts.
To further explore the impact of target shift on PtO similarity, we analyze the Warcraft setting under two PtO tasks: minimizing path cost alone and minimizing both path cost and length. Using the same experimental procedure as in Section 5.2, we apply it to these tasks while considering target shifts in feature-label datasets. Although the same target shifts are applied, their effect on PtO transferability is less severe for minimizing both path cost and length compared to minimizing cost alone (Fig. 5). Our decision-aware dataset distance, using weights from Section 5.2, effectively captures this behavior. The distance distribution for the task less impacted by the target shift is more left-skewed (Fig. 5(b)). In contrast, the dataset distance that only accounts for features and labels, is unable to differentiate between these two tasks (Fig. 5(a)).
(a) Not including decisions
(b) Including decisions
Figure 5: Distance vs. Adaptation for two tasks in the Warcraft setting. Dataset distance is computed (a) without incorporating decisions, and (b) with decision incorporation.
6 Discussion
In this work, we introduce the first notion of dataset distances specifically tailored to PtO tasks. Our decision-aware dataset distance integrates features, labels, and decisions, enhancing the evaluation of prediction-to-decision similarities across tasks. Experiment results show that including the decision component significantly boosts transferability, particularly in complex environments where label changes do not clearly correlate with decisions. This integration allows our metric to effectively capture nuances dictated by the structure of downstream optimization problems, demonstrating robust reflection of task dynamics without direct analysis of these structures.
Our framework is adaptable and can handle diverse PtO tasks, allowing for flexible weighting of each component to provide meaningful comparisons across a broad range of applications. This adaptability is crucial for realistic applications where datasets vary not only in features and labels but also in the nature of the decisions they inform.
Looking forward, there are several promising avenues for expanding our framework. Extending our framework to handle decision components of varying structures and dimensions using techniques like the Gromov-Wasserstein distance or approaches such as those in Alvarez-Melis and Fusi (2020) for comparing unrelated labels, could bridge gaps between non-comparable decision spaces. Further refining the weights assigned to features, labels, and decisions, and exploring tuning methods independent of transferability measures, could greatly improve the utility of our metric, especially in unlabeled scenarios. Additionally, adapting our approach to more intricate PtO structures—such as when multiple feature-label pairs define a single decision— through a hierarchical OT framework (Yurochkin et al., 2019) could extend its applicability, enhancing the robustness and versatility of our method across various PtO scenarios. By establishing this initial notion of dataset distance designed for PtO tasks, our work provides a foundational step for future research in this area.
Acknowledgments
This material is based upon work supported by the AI Research Institutes Program funded by the National Science Foundation under AI Institute for Societal Decision Making (AI-SDM), Award No. 2229881.
Appendix A Our decision-aware distance is a well defined metric
Proposition A.1.
For any $\bm{\alpha}=(\alpha_{X},\alpha_{Y},\alpha_{W})$ with $\alpha_{X},\alpha_{Y},\alpha_{W}>0$, the decision-aware dataset distance $d_{OT}(\mathcal{D},\mathcal{D}^{\prime};c_{PtO}^{\bm{\alpha}})$ defines a valid metric on $\mathcal{P}(\mathcal{X}\times\mathcal{Y}\times\Omega)$, where $\mathcal{P}(\mathcal{X}\times\mathcal{Y}\times\Omega)$ represents the space of probability measures over joint distributions of features $\mathcal{X}$, labels $\mathcal{Y}$, and decisions $\Omega$. If $\alpha_{Y}=0$, then the decision-aware dataset distance is still at least a pseudometric.
Proof.
To demonstrate that $d_{OT}(\cdot,\cdot;c_{PtO}^{\mathbf{\alpha}})$ is a valid metric, it is sufficient to verify that the ground cost function $c_{PtO}$ used in the optimal transport problem is a metric on $\mathcal{X}\times\mathcal{Y}\times\Omega$. If $c_{PtO}$ is indeed a metric, then $d_{OT}(\cdot,\cdot;c_{PtO})$ corresponds to the Wasserstein distance [Villani, 2008]. In Equation 4.1, $d_{OT}(\cdot,\cdot;c_{PtO})$ is defined as a convex combination of $d_{\mathcal{X}}$ and $d_{\mathcal{Y}}$, which are metrics on $\mathcal{X}$ and $\mathcal{Y}$ respectively, and the decision quality disparity $l_{q}$. To show that $c_{PtO}$ is a metric, it suffices to show that $l_{q}$ satisfies the four metric properties: non-negativity, identity of indiscernibles, symmetry, and the triangle inequality. If $l_{q}$ does not individually satisfy these properties, we must demonstrate that the convex combination of $d_{\mathcal{X}}$, $d_{\mathcal{Y}}$, and $l_{q}$ satisfies these properties collectively under the assumption that $\alpha_{X},\alpha_{Y},\alpha_{W}>0$.
First, $l_{q}$ is clearly non-negative because it is defined as an absolute value. It is symmetric in the convex combination of $c_{PtO}$ because it is taken as the absolute difference between two decision qualities with fixed true costs.
$l_{g}(z,z^{\prime};y^{\prime},y^{\prime})$ $=\big{|}q(z;y^{\prime})-q(z^{\prime};z^{\prime})\big{|}$
$=\big{|}q(z^{\prime};y^{\prime})-q(z;z^{\prime})\big{|}$
$=l_{g}(z^{\prime},z;y^{\prime},y^{\prime})$
Moreover, $l_{q}$ satisfies triangle inequality due to the triangle inequality property of the absolute value.
$l_{g}(z_{1},z_{2};y_{1},y_{2})+l_{g}(z_{2},z_{3};y_{2},y_{3})$
$=\big{|}g(z_{1};y_{1})-g(z_{2};y_{2})\big{|}+\big{|}g(z_{2};y_{2})-g(z_{3};y_{3})\big{|}$
$\leq\big{|}g(z_{1};y_{1})-g(z_{2};y_{2})+g(z_{2};y_{2})-g(z_{3};y_{3})\big{|}$
$=\big{|}g(z_{1};y_{1})-g(z_{3};y_{3})\big{|}$
$=l_{g}(z_{1},z_{3};y_{1},y_{3})$
Lastly, while $l_{q}$ might not satisfy the identity of indiscernibles in isolation (specifically, $l_{q}(y,y^{\prime};z,z)=0$ does not necessarily imply $y=y^{\prime}$; meaning two different decisions can lead to the same objective value), $c_{PtO}$ does satisfy this property for $\alpha_{X},\alpha_{Y},\alpha_{W}>0$. If $(x,y,z)=(x^{\prime},y^{\prime},z^{\prime})$, then $l_{g}(z,z^{\prime};y^{\prime},y^{\prime})=\big{|}g(z;y)-g(z^{\prime};y)\big{|}=0$ because $z=z^{\prime}$ implies $g(z;y)=g(z^{\prime};y)$ and hence $c_{PtO}((x,y,z),(x^{\prime},y^{\prime},z^{\prime}))=0$. Conversely, if $c_{PtO}((x,y,z),(x^{\prime},y^{\prime},z^{\prime}))=0$, then $d_{\mathcal{X}}(x,x^{\prime})=0$, $d_{\mathcal{Y}}(y,y^{\prime})=0$, and $l_{q}(y,y^{\prime};z,z)=0$ because $\alpha_{X},\alpha_{Y},\alpha_{W}>0$. Since $d_{\mathcal{Y}}(y,y^{\prime})=0$ implies $y=y^{\prime}$ (because $d_{\mathcal{Y}}$ is a metric), it follows that $w^{*}(y)=w^{*}(y^{\prime})$ and hence $z=z^{\prime}$.
Therefore, $c_{PtO}$ satisfies the identity of indiscernibles. Consequently, since $l_{q}$ satisfies non-negativity, symmetry, and the triangle inequality, and since $c_{PtO}$ satisfies the identity of indiscernibles, $d_{OT}(\cdot,\cdot;c_{PtO})$ is indeed a valid metric with $c_{PtO}$ a valid metric on $\mathcal{X}\times\mathcal{Y}\times\Omega$. ∎
Appendix B Implications of Assumption 4.3
Assumption 4.3
The decision quality function is $k_{1},k_{2}$-Lipschitz. This means that for any $y,y^{*},z,z^{*}\in\mathcal{Y}$ the following inequality holds:
$|q(y,y^{*})-q(z,z^{*})|\leq k_{1}\norm{y-z}+k_{2}\norm{y^{*}-z^{*}}.$
To establish the bound presented in Theorem 4.4, we rely on the fact that $l_{g}$ is $k_{1},k_{2}$-Lipschitz under Assumption 4.3. The following proposition demonstrates that $l_{g}$ indeed satisfies the Lipschitz condition given this assumption.
Proposition B.1.
If $g$, the objective function of the downstream optimization problem, is $k_{1},k_{2}$-Lipschitz (Assumption 4.3), then $l_{g}$ is also $k_{1},k_{2}$-Lipschitz.
Proof.
$\big{|}l_{g}(z,z_{1};y,y_{1})-l_{g}(z,z_{2};y,y_{2})\big{|}$
=|$=\big{|}\absolutevalue{g(z;y)-g(z_{1};y_{1})}-\absolutevalue{g(z;y)-g(z_{2};y_{2})}\big{|}$ | |
$\leq\big{|}g(z;y)-g(z_{1};y_{1})-g(z;y)+g(z_{2};y_{2})\big{|}$ (8)
$=\big{|}g(z_{2};y_{2})-g(z_{1};y_{1})\big{|}$
$=\big{|}g(z_{2};y_{2})-g(z_{1};y_{2})+g(z_{1};y_{2})-g(z_{1};y_{1})\big{|}$
$\leq\big{|}g(z_{2};y_{2})-g(z_{1};y_{2})\big{|}+\big{|}g(z_{1};y_{2})-g(z_{1};y_{1})\big{|}$ (9)
$\leq k_{1}\norm{z_{1}-z_{2}}+k_{2}\norm{y_{1}-y_{2}}$ (10)
Inequalities (8) and (9) are a result of the triangle inequality of the absolute value. Inequality (10) is due to the $k_{1}-k_{2}$-lipschitzness of $g$. ∎
Assumption 4.3 imposes a specific structure on the downstream optimization problem by assuming that the decision quality function has a bounded rate of change with respect to both the predicted and true cost vectors. This is a reasonable assumption for certain downstream optimization tasks, as highlighted in the following lemmas.
Lemma B.2.
If $M(\cdot)$ is a convex program with a strongly convex objective and constraints with independent derivatives (Linear Independence Constraint Qualification), Assumption 4.3 holds.
Lemma B.3.
If $M(\cdot)$ has a linear optimization objective with a strongly convex feasible region, Assumption 4.3 holds.
Appendix C Proof of Theorem 4.4
Definition C.1 (Probabilistic Transfer Lipschitzness [Courty et al., 2017]).
Let $\mu_{1}$ and $\mu_{2}$ be distributions over some metric space $\mathcal{X}$ with metric $d_{\mathcal{X}}$. Let $\Pi(\mu_{1},\mu_{2})$ be a joint distribution (also referred to as coupling) over $\mu_{1}\times\mu_{2}$. Let $\phi:\mathbb{R}\to[0,1]$. A labeling function $f:\mathcal{X}\to\mathbb{R}$ is $\phi$-Lipschitz transferable with respect to $\Pi$ if for all $\lambda>0$: $\Pr_{(x_{1},x_{2})\sim\Pi(\mu_{1},\mu_{2})}[|f(x_{1})-f(x_{2})|>\lambda d_{\mathcal{X}}(x_{1},x_{2})]\leq\phi(\lambda)$
Theorem C.2.
Suppose Assumption 4.3 holds. For a feature space $\mathcal{X}$, a label space $\mathcal{Y}$, and a decision set $\Omega$, let $\mathcal{W}:=\mathcal{X}\times\mathcal{Y}\times\Omega$. Let $\mathcal{P}_{S}$ and $\mathcal{P}_{T}$ be the source and target distributions over $\mathcal{X}\times\mathcal{Y}$ respectively. For any labeling function $f:\mathcal{X}\to\mathcal{Y}$, let $\mathcal{P}_{T}^{f}$ and $\mathcal{P}_{S}^{*}$ be distributions over $\mathcal{W}$ given by $\mathcal{P}_{T}^{f}:=(x,y,w^{*}(f(x)))_{(x,y)\sim\mathcal{P}_{T}}$ and $\mathcal{P}_{S}^{*}:=(x,y,w^{*}(y))_{(x,y)\sim\mathcal{P}_{S}}$. For a ground cost function of the form
$c_{PtO}^{\bm{\alpha}}((x,y,z),(x^{\prime},y^{\prime},z^{\prime}))=\alpha_{X}d_{\mathcal{X}}(x,x^{\prime})+\alpha_{Y}d_{\mathcal{Y}}(y,y^{\prime})+\alpha_{W}l_{g}(z,z^{\prime};y^{\prime},y^{\prime}),$
let $\Pi^{*}$ be the coupling that minimizes the OT problem with ground cost $c_{PtO}^{\bm{\alpha}}$ between $\mathcal{P}_{T}^{f}$ and $\mathcal{P}_{S}^{*}$. Let $\tilde{f}$ be a labeling function that is $\phi$-Lipschitz transferable w.r.t. $\Pi^{*}$. We assume $\mathcal{X}$ is bounded by $K$ and $\tilde{f}$ is $l$-Lipschitz, such that $|\tilde{f}(x_{1})-\tilde{f}(x_{2})|\leq 2lK=L$. Then, for all $\lambda>0$ and $\alpha_{W}\in(0,1)$ such that $(\lambda k_{1}+k_{2}+1)\alpha_{W}=1$, and $\alpha_{X}=\lambda k_{1}\alpha_{W}$ and $\alpha_{Y}=k_{2}\alpha_{W}$, we have with probability at least $1-\delta$ that:
$err(f;q_{\text{reg}},\mathcal{P}_{T})\leq\ err(\tilde{f};q_{\text{reg}},\mathcal{P}_{S})+err(\tilde{f};q_{\text{reg}},\mathcal{P}_{T})+k_{1}L\phi(\lambda)+(1/\alpha_{W})d_{OT}(\mathcal{P}_{T}^{f},\mathcal{P}_{S}^{*}\ ;c_{PtO}^{\bm{\alpha}})$
Proof.
$er$ $r(f;q_{\text{reg}},\mathcal{P}_{T})$
$=\mathbb{E}_{(x,y)\sim\mathcal{P}_{T}}\ l_{g}(w^{*}(f(x)),w^{*}(y);y,y)$
$\leq\mathbb{E}_{(x,y)\sim\mathcal{P}_{T}}\ l_{g}(w^{*}(f(x)),w^{*}(\tilde{f}(x));y,y)+\mathbb{E}_{(x,y)\sim\mathcal{P}_{T}}\ l_{g}(w^{*}(\tilde{f}(x)),w^{*}(y);y,y)$ (11)
$=\mathbb{E}_{(x,y)\sim\mathcal{P}_{T}}\ l_{g}(w^{*}(\tilde{f}(x)),w^{*}(f(x));y,y)+\mathbb{E}_{(x,y)\sim\mathcal{P}_{T}}\ l_{g}(w^{*}(\tilde{f}(x)),w^{*}(y);y,y)$ (12)
$=\mathbb{E}_{(x,y,z)\sim\mathcal{P}_{T}^{f}}\ l_{g}(w^{*}(\tilde{f}(x)),z;y,y)+\mathbb{E}_{(x,y)\sim\mathcal{P}_{T}}\ l_{g}(w^{*}(\tilde{f}(x)),w^{*}(y);y,y)$ (13)
$=\mathbb{E}_{(x,y,z)\sim\mathcal{P}_{T}^{f}}\ l_{g}(w^{*}(\tilde{f}(x)),z;y,y)-err(\tilde{f};q_{\text{reg}},\mathcal{P}_{S})+err(\tilde{f};q_{\text{reg}},\mathcal{P}_{S})+err(\tilde{f};q_{\text{reg}},\mathcal{P}_{T})$
$=\mathbb{E}_{(x,y,z)\sim\mathcal{P}_{T}^{f}}\ l_{g}(w^{*}(\tilde{f}(x)),z;y,y)-\mathbb{E}_{(x,y,z)\sim\mathcal{P}_{S}^{*}}\ l_{g}(w^{*}(\tilde{f}(x)),z;y,y)+err(\tilde{f};q_{\text{reg}},\mathcal{P}_{S})+err(\tilde{f};q_{\text{reg}},\mathcal{P}_{T})$
$\leq\big{|}\mathbb{E}_{(x,y,z)\sim\mathcal{P}_{T}^{f}}\ l_{g}(w^{*}(\tilde{f}(x)),z;y,y)-\mathbb{E}_{(x,y,z)\sim\mathcal{P}_{S}^{*}}\ l_{g}(w^{*}(\tilde{f}(x)),z;y,y)\big{|}+err(\tilde{f};q_{\text{reg}},\mathcal{P}_{S})+err(\tilde{f};q_{\text{reg}},\mathcal{P}_{T})$
Inequality (11) uses the fact that $l_{g}(\ \cdot\ ;y,y)$ satisfies the triangle inequality and line (12) is due to the symmetry of $l_{g}(\ \cdot\ ;y,y)$ for any $y\in\mathcal{Y}$. Line (13) comes from the fact that $\mathcal{P}_{T}^{f}:=(x,f(x),y)_{(x,y)\sim\mathcal{P}_{T}}$. We continue by bounding the first term. {adjustwidth}-0.4cm0pt
$\big{|}\mathbb{E}_{(x,y,z)\sim\mathcal{P}_{T}^{f}}\ l_{g}(w^{*}(\tilde{f}(x)),z;y,y)-\mathbb{E}_{(x,y,z)\sim\mathcal{P}_{S}^{*}}\ l_{g}(w^{*}(\tilde{f}(x)),z;y,y)\big{|}$
=|$=\absolutevalue{\int_{\mathcal{W}}l_{g}(w^{*}(\tilde{f}(x)),z;y,y)(\mathcal{P}_{T}^{f}(X=x,Y=y,Z=z)-\mathcal{P}_{S}^{*}(X=x,Y=y,Z=z))\mathop{}\!\mathrm{d}x\mathop{}\!\mathrm{d}y\mathop{}\!\mathrm{d}z}$ |
=|$=\absolutevalue{\int_{\mathcal{W}}l_{g}(w^{*}(\tilde{f}(x)),z;y,y)\mathop{}\!\mathrm{d}\Pi^{*}((x_{s},y_{s},z_{s}),(x_{t},y_{t},z_{t}^{f}))}$ |
$\leq\int_{\mathcal{W}^{2}}\absolutevalue{l_{g}(\tilde{z}_{t},z_{t}^{f};y_{t},y_{t})-l_{g}(\tilde{z}_{s},z_{s};y_{s},y_{s})}\mathop{}\!\mathrm{d}\Pi^{*}(\mathbf{w}_{s},\mathbf{w}_{t}^{f})$ (14)
$\leq\int_{\mathcal{W}^{2}}\Big{|}l_{g}(\tilde{z}_{t},z_{t}^{f};y_{t},y_{t})-l_{g}(\tilde{z}_{s},z_{t}^{f};y_{s},y_{t})\Big{|}+\Big{|}l_{g}(\tilde{z}_{s},z_{t}^{f};y_{s},y_{t})-l_{g}(\tilde{z}_{s},z_{s},;y_{s},y_{s})\Big{|}\mathop{}\!\mathrm{d}\Pi^{*}(\mathbf{w}_{s},\mathbf{w}_{t}^{f})$ (15)
$\leq\int_{\mathcal{W}^{2}}k_{1}d_{\mathcal{Y}}(\tilde{f}(x_{t}),\tilde{f}(x_{s}))+k_{2}d_{\mathcal{Y}}({y_{t},y_{s}})+\Big{|}l_{g}(\tilde{z}_{s},z_{t}^{f};y_{s},y_{t})-l_{g}(\tilde{z}_{s},y_{s};y_{s},y_{s})\Big{|}\mathop{}\!\mathrm{d}\Pi^{*}(\mathbf{w}_{s},\mathbf{w}_{t}^{f})$ (16)
$\leq\ k_{1}L\phi(\lambda)+\int_{\mathcal{W}^{2}}\lambda k_{1}d_{\mathcal{X}}(x_{t},x_{s})+k_{2}d_{\mathcal{Y}}({y_{t},y_{s}})+\Big{|}l_{g}(\tilde{z}_{s},z_{t}^{f};y_{s},y_{t})-l_{g}(\tilde{z}_{s},y_{s};y_{s},y_{s})\Big{|}\mathop{}\!\mathrm{d}\Pi^{*}(\mathbf{w}_{s},\mathbf{w}_{t}^{f})$ (17)
$\leq\ k_{1}L\phi(\lambda)+\int_{\mathcal{W}^{2}}\lambda k_{1}d_{\mathcal{X}}(x_{t},x_{s})+k_{2}d_{\mathcal{Y}}({y_{t},y_{s}})+l_{g}(z_{t}^{f},z_{s};y_{s},y_{s})\mathop{}\!\mathrm{d}\Pi^{*}(\mathbf{w}_{s},\mathbf{w}_{t}^{f})$
From line (14) onwards we take $\mathbf{w}_{s}:=(x_{s},y_{s},y_{s}),\mathbf{w}_{t}^{f}:=(x_{t},y_{t}^{f},y_{t})$ and $\tilde{z}_{s}=w^{*}(\tilde{f}(x_{s})),\tilde{z}_{t}=w^{*}(\tilde{f}(x_{t}))$ for ease of notation. Given a weight $\alpha_{W}$, we now normalize the last term such that the ground cost function is a convex combination of $d_{\mathcal{X}}$, $d_{\mathcal{Y}}$m and $l_{g}$.
$\int_{\mathcal{W}^{2}}\lambda k_{1}d_{\mathcal{X}}(x_{t},x_{s})+k_{2}d_{\mathcal{Y}}({y_{t},y_{s}})+l_{g}(z_{t}^{f},z_{s};y_{s},y_{s})\mathop{}\!\mathrm{d}\Pi^{*}(\mathbf{w}_{s},\mathbf{w}_{t}^{f})$
$=\frac{1}{\alpha_{W}}\int_{\mathcal{W}^{2}}\lambda k_{1}\alpha_{W}d_{\mathcal{X}}(x_{t},x_{s})+k_{2}\alpha_{W}d_{\mathcal{Y}}(x_{t},x_{s})+\alpha_{W}l_{g}(z_{t}^{f},z_{s};y_{s},y_{s})\mathop{}\!\mathrm{d}\Pi^{*}(\mathbf{w}_{s},\mathbf{w}_{t}^{f})$
$=\frac{1}{\alpha_{W}}\ d_{OT}(\mathcal{P}^{f}_{T},\mathcal{P}^{*}_{S};c_{PtO}^{\bm{\alpha}})$
Appendix D Details on Experimental Settings
D.1 Linear Model Top-K
D.2 Problem setting
The Linear Model Top-K setting, first introduced by Shah et al. [2022], is a learning task within the Predict-then-Optimize (PtO) framework. The objective is to train a linear model to perform top $K$ selection on data that is inherently generated by a cubic polynomial function. This setting provides a controlled environment to assess the effectiveness of decision-focused learning methods when the underlying relationship between features and outcomes is nonlinear, yet the model used for prediction is linear. The PtO task in this setting is as follows:
  • Predict: Given the feature $x_{n}\sim\mathcal{P}_{\mathcal{X}}$, where $\mathcal{P}_{\mathcal{X}}=\text{Unif}[-1,1]$, of a resource $n$, the prediction tasks consists of using a linear model to predict the corresponding utility $\hat{y}_{n}$, where the true utility $y_{n}=p(x_{n})$ is a cubic polynomial in $x_{n}$. The predictions for $N$ resources are aggregated into a vector $\hat{\bm{y}}=[\hat{y}_{1},\ldots,\hat{y}_{N}]$, where each element corresponds to the predicted utility of a resource.
  • Optimize: The optimization task involves selecting the $K$ out of $N$ resources with the highest utility. This corresponds to solving the optimization problem $M(\hat{\bm{y}})=\max_{\bm{z}\in[0,1]^{N}}\{\bm{z}\cdot\sigma_{x}(\hat{\bm{y}})\}$ such that $||\bm{z}||_{0}=K$, where $\sigma_{x}$ is the permutation that orders $\hat{\bm{y}}$ in ascending order of $\bm{x}=[x_{1},\ldots,x_{N}]$.
D.3 Distribution shift
We introduce synthetic distribution shifts by modifying the original feature-label distribution $\mathcal{P}=(\text{Id},p)_{*}U[-1,1]$. Specifically, for various values of $\gamma\in[0,1.3]$, we define the feature-label distributions $\mathcal{P}_{\gamma}=(\text{Id},p_{\gamma})_{*}U[-1,1]$ where $p_{\gamma}(x_{n})=10(x_{n}^{3}-\gamma x_{n})$ (illustrated in Fig. 6), using $\mathcal{P}_{0.65}$ as the target distribution.
D.4 Model training
To train models in the target and source datasets we use a surrogate loss function as done in Shah et al. [2022]. Specifically, we use the entropy regularized Top-K loss function proposed by Xie et al. [2020] that reframes the Top-K problem with entropy regularization as an optimal transport problem.
Figure 6: Distribution shifts considered for source datasets in the Linear Model TopK Setting.
D.5 Warcraft Shortest Path
D.5.1 Problem setting
Adapted from Mandi et al. [2023] and Tang and Khalil [2023], this setting involves finding the minimum cost path on $d\times d$ RGB grid maps from the Warcraft II tileset dataset [Vlastelica et al., 2020], where each pixel has an unknown cost. The task is to predict these costs and then determine the minimum cost path from the top-left to the bottom-right pixel.
  • Predict: Given the feature $x_{n}\in\mathbb{R}^{d\times d}$, predict traveling costs $\hat{y}_{n}$ as a grid in $\mathbb{R}^{p\times p}$ with $d\geq p$.
  • Optimize: Find the path with the minimum traveling cost from the left top corner to the bottom right corner of $\hat{y}_{n}$. This corresponds to solving the optimization problem $M(\hat{\bm{y}})=\min_{\bm{z}\in[0,1]^{p}}\{\bm{z}\cdot\hat{y}\}$ such that $\bm{z}_{0,0}=\bm{z}_{p,p}=1$ and $\bm{z}$ represents a connected path.
D.5.2 Distribution shift
The original distribution $\mathcal{P}$, which we treat as the target distribution, is defined over $\mathbb{R}^{d\times d}\times\mathbb{R}^{p\times p}$, where $d=96$ and $p=12$. Here, $\mathbb{R}^{d\times d}$ represents the feature space depicting maps, while $\mathbb{R}^{p\times p}$ represents the traveling costs on these maps. We induce a target shift for $\mathcal{P}_{\gamma}$ by uniformly sampling the costs for different pixel classes from the same range as $\mathcal{P}$ ([0.8, 9.2] for the Warcraft II tileset dataset). Figure 7 illustrates the costs coming from two different distributions over one same feature while highlighting the different decisions (shortest path) that these costs yield.
Figure 7: Synthetic distribution shift in Warcraft Shortest Path. The white line illustrates the decision, corresponding to the shortest path, on dataset A (center) and dataset B (right) for a sample with the same features (left map).
D.6 Inventory Stock Problem
D.6.1 Problem setting
In this problem, adapted from Donti et al. [2017], a company must determine the optimal order quantity $z$ of a product to minimize costs given a stochastic demand $y$, which is influenced by observed features $x$. The cost structure includes both linear and quadratic costs for the amount of product ordered, as well as different linear and quadratic costs for over-orders $[z-y]^{+}$ and under-orders $[y-z]^{+}$. The objective function is:
$f_{\text{stock}}(y,z)=$ $c_{0}z+\frac{1}{2}q_{0}z^{2}+c_{b}[y-z]_{+}+\frac{1}{2}q_{b}([y-z]_{+})^{2}$
$+c_{h}[z-y]_{+}+\frac{1}{2}q_{h}([z-y]_{+})^{2}$ (18)
[…]
To simplify the setting, we assume that the demands are discrete, taking on values $d_{1},\ldots,d_{k}$ with probabilities (conditional on $x$) $\left(p_{\theta}\right)_{i}\equiv p\left(y=d_{i}|x;\theta\right)$. Thus, our stochastic programming problem can be succinctly expressed as a joint quadratic program:
$\underset{z\in\mathbb{R},z_{b},z_{h}\in\mathbb{R}^{k}}{\operatorname{minimize}}\Big{\{}c_{0}z+\frac{1}{2}q_{0}z^{2}+\sum_{i=1}^{k}\left(p_{\theta}\right)_{i}\big{(}c_{b}(z_{b})_{i}$ (10)
$\hskip 62.59596pt+\frac{1}{2}q_{b}(z_{b})_{i}^{2}+c_{h}(z_{h})_{i}+\frac{1}{2}q_{h}(z_{h})_{i}^{2}\big{)}\Big{\}}$
$\text{subject to}\quad d-z\mathbf{1}\leq z_{b},\quad z\mathbf{1}-d\leq z_{h},\quad z,z_{h},z_{b}\geq 0$
D.6.2 Distribution shift
We generate problem instances by randomly sampling $x\in\mathbb{R}^{n}$ and then generating $p(y|x;\theta)$ according to $p(y|x;\theta)\propto\exp((\Theta^{T}x)^{2})$. We introduce distribution shifts for both $x$ and $y$. Specifically, $x$ is sampled from a Gaussian distribution where the mean is sampled from $U[-0.5,0.5]$, and $\Theta$ is also sampled from a Gaussian distribution.
Appendix E Additional Results
In Section 5.2 we analyzed the correlation between dataset distance and transferability in PtO. The plots presented in Figure 8 show this correlation for the Linear Model TopK setting and the Inventory Stock problem for two weighting profiles: the best weighting profile where no decisions are included [left] and the best weighting profile with decisions included [right]. For both setting, decisions improve predictability of transferability. This improvement is less pronounced in the Inventory Stock problem.
Figure 8: Distance vs Adaptation. OT distance for the best feature-label and feature-label-decision weighting against the regret transferability.